Find m∠RSQ and m∠TSQ. S T Q R (8x + 18)° (15x − 43)

RSQ= __° 

TSQ= __°


1
Expert's answer
2020-09-14T18:46:37-0400

they add to 90"\degree" so

∠TSQ + ∠RSQ = 90"\degree"

(15x - 43)"\degree" + (8x + 18)"\degree" = 90"\degree"

add like terms

23x"\degree" - 25"\degree" = 90"\degree"

add 25 to bothn sides

23x"\degree" = 115"\degree"

divide both sides by 23"\degree"

"\frac{23x\degree}{23\degree}" = "\frac{115\degree}{23\degree}"


x = 5"\degree"


NOW ∠RSQ = (8x + 18)"\degree"

= {8(5) + 18}"\degree"

= ( 40 + 18 )"\degree"

∠RSQ = 58"\degree"

the other angle is just

90"\degree" - ∠RSQ = ∠TSQ

90"\degree" - 58"\degree" = ∠TSQ

32"\degree" = ∠TSQ


SO, ∠TSQ = 32"\degree"

∠RSQ = 58"\degree"